Circular No. 6904

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998bx, 1998by, 1998bz, 1998ca, 1998cb O. Perdereau and N. Regnault, Laboratoire de l’Accélérateur Linéaire,

Université de Paris-Sud, on behalf of the EROS collaboration, report the discovery of five apparent supernovae, discovered in the course of the automated supernova search with the 1-m Marly telescope (+ CCD) at La Silla (limiting mag ∼ 20.2). Finding charts can be found at http://www.lal.

in2p3.fr/EROS/SNe/Charts/.

SN  1997 UT  α_{2000}  δ_{2000}  V  Offset
1998bx Apr. 21 11^{h}12^{m}28.^{s}05 − 5^{o}58′09′′.9 19.6 1′′.2 W, 5′′ N
1998by Apr. 21 11 25 07.20 − 5 23 15.2 19.5 0′′.6 E, 0′′.6 S
1998bz Apr. 24 12 11 05.84 − 5 35 33.8 19.1 near center
1998ca Apr. 24 12 13 29.24 − 6 23 42.0 18.8 7′′ W, 7′′ N
1998cb Apr. 28 10 40 16.07 − 6 15 17.2 20.0 2′′.3 W, 0′′.8 N

SN 1998bx was confirmed on CCD frames taken on Apr. 22, but was not present on a frame taken on 1997 Mar. 7. SN 1998by was confirmed on CCD frames taken on Apr. 22, but was not present on a frame taken on 1997 Mar. 15. SN 1998bz and SN 1998ca were found by differencing CCD frames taken on 1997 Mar. 7 and 1998 Apr. 24, and were confirmed on CCD frames taken on Apr. 29. SN 1998cb was confirmed on CCD frames taken on Apr. 29, but was not present on a frame taken on 1997 Apr. 29.

XTE J1855–026 F. E. Marshall, Goddard Space Flight Center (GSFC); and T. Take-

shima, Universities Space Research Association and GSFC, report the detection of a transient x-ray pulsar XTE J1855–026 with the Proportional Counter Array on the Rossi X-ray Timing Explorer: “The peak observed 2–10-keV flux was 13 mCrab on May 6.75 UT. A pointed observation on May 12.85, when the flux had faded to 6 mCrab, revealed periodic modulation of the flux with a peak-to-peak amplitude of ∼ 60 percent and a period of 362 ± 1 s. The best-fit position is α = 18^{h}55^{m}.7, δ = −2^{o}37′ (equinox 2000.0; estimated uncertainty 6′). The source is probably a Be/neutron star binary system. We note that the error circle contains the Be star BD −02^{o}4786. Optical observations are strongly encouraged.”

SUPERNOVA 1998bv IN HS 1035+4758

Corrigendum. On IAUC 6900, line 1, Nizhnij Arkhyz is a site, not a person.

1998 May 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green